. The members of a flying club plan to share
equally the cost of a $200,000 airplane. The members
want to find five more people to join the club so that the
cost per person will decrease by $2000. How many members
are currently in the club?

Answer by stanbon(75887)

The members of a flying club plan to share
equally the cost of a $200,000 airplane. The members
want to find five more people to join the club so that the
cost per person will decrease by $2000. How many members
are currently in the club?
----------------
Original # of members is "x".
Average cost = 200,000/x
-------
New # of members is "x+5".
Average cost = 200,000,(x+5).
---
Equation:
old average - new average = 2000
200,000/x - 200,000/(x+5) = 2000
---
Divide thru by 2000 to get:
100/x - 100/(x+5) = 1
Multiply thru by x(x+5) to get:
100(x+5) - 100x = x (x+5)
500 = x^2+5x
---
x^2+5x-500 = 0
(x+25)(x-20) = 0
---
Positive solution:
x = 20 (original # of members)
===================================
Cheers,
Stan H.
============